THE MAGELLANIC CLOUDS NEWSLETTER

An electronic publication dedicated to the Magellanic Clouds, and astrophysical phenomena therein

No. 108 — 5 December 2010

http://www.astro.keele.ac.uk/MCnews

Editor: Jacco van Loon

Editorial

It is my pleasure to present you the 108^{th} issue of the Magellanic Clouds Newsletter. Slightly delayed, in part because of an observing run of the Magellanic Clouds... uhum, well... we did not get to point the telescope at them but the night before we were meant to start a PhD student did see them for the first time, and she was suitably in awe. That’s what we do it for!

This issue contains no fewer than five very interesting papers regarding the origin of the Magellanic Clouds and their interaction: Růžička, Theis & Palouš; Boylan-Kolchin, Besla & Hernquist; Yang & Hammer; Liu et al.; Busha et al. Another, highly original piece of work concerns the experiment to measure the interstellar scintillation of Magellanic Cloud stars as their light passes through the Milky Way’s turbulent interstellar medium (Habibi et al.).

There are exciting jobs out there, at the Kavli institute in fast-developing China, or at Macquarie University to use the successor and surpasser of 2dF/AAΩ.

The next issue is planned to be distributed on the 1^{st} of February 2011. With the best wishes for the New Year (on some calendars),

Editorially Yours, Jacco van Loon

Refereed Journal Papers

The kinematics and chemical stratification of the type Ia supernova remnant 0519−69.0. An XMM–Newton and Chandra study

D. Kosenko^{1,2}, E.A. Helder^{1} and J. Vink^{1}

1Astronomical Institute Utrecht, Utrecht University, P.O. Box 80000, 3508TA Utrecht, The Netherlands 2Sternberg Astronomical Institute, 119992, Universitetski pr., 13, Moscow, Russia

We present a detailed analysis of the XMM–Newton and Chandra X-ray data of the young type Ia supernova remnant SNR 0519−69.0, which is situated in the Large Magellanic Cloud. We used data from both the Chandra ACIS and XMM–Newton EPIC MOS instruments, and high resolution X-ray spectra obtained with the XMM–Newton reflection grating spectrometer (RGS). Our analysis of the spatial distribution of X-ray line emission using the Chandra data shows that there is a radial stratification of oxygen, intermediate mass elements (IME) and iron, with the emission from more massive elements peaking more toward the center. Using a deprojection technique we measure a forward shock radius of 4.0±0.3 pc and a reverse shock radius of 2.7±0.4 pc. We took the observed stratification of the shocked ejecta into account in the modeling of the X-ray spectra, for which we used multi-component non-equilibrium ionization models, with the components corresponding to layers dominated by one or two elements. An additional component was added in order to represent the shocked interstellar medium, which mostly contributed to the continuum emission. This multicomponent model fits the data adequately, and was also employed to characterize the spectra of distinct regions extracted from the Chandra data. From our spectral analysis we find that the approximate fractional masses of shocked ejecta for the most abundant elements are: M_{O} = 32%, M_{Si/S} = 7%/5%, M_{Ar+Ca} = 1% and M_{Fe} = 55%. From the continuum component we derive a circumstellar density of n_{H} = 2.4 ± 0.2 cm^{−3}. This density, together with the measurements of the forward and reverse shock radii suggest an age of 0519−69.0 of 450 ± 200 yr, somewhat lower than, but consistent with the age estimate based on the extent of the light echo (600 ± 200 yr). Finally, from the high resolution RGS spectra we measured a Doppler broadening of σ = 1873 ± 50 km s^{−1}, from which we derive a forward shock velocity of v_{FS} = 2770±500 km s^{−1}. We discuss our results in the context of single degenerate explosion models, using semi-analytical and numerical modeling, and compare the characteristics of 0519−69.0 with those of other type Ia supernova remnants.

Published in Astronomy and Astrophysics, 519A, 11 (2010) Available from arXiv:1001.0983

Anchoring the universal distance scale via a Wesenheit template

Daniel J. Majaess^{1,2}, David G. Turner^{1,2}, David J. Lane^{1,2}, Arne Henden^{3,4,5} and Tom Krajci^{3,4}

1Saint Mary’s University, Halifax, Nova Scotia, Canada 2Abbey Ridge Observatory, Stillwater Lake, Nova Scotia, Canada 3American Association of Variable Star Observers, Cambridge, MA, USA 4Astrokolkhoz Telescope Facility, Cloudcroft, New Mexico, USA 5Sonoita Research Observatory, Sonoita, Arizona, USA

A V I Wesenheit diagram featuring SX Phoenicis, δ Scuti, RR Lyrae, type II and classical Cepheid variables is calibrated by means of geometric-based distances inferred from HST, Hipparcos, and VLBA observations (n = 30). The distance to a target population follows from the offset between the observed Wesenheit magnitudes and the calibrated template. The method is evaluated by ascertaining the distance moduli for the LMC (μ_{0} = 18.43 ± 0.03 (σx̄) and the globular clusters ω Cen, M 54, M 13, M 3, and M 15. The results agree with estimates cited in the literature, although a nearer distance to M 13 is favoured (pending confirmation of the data’s photometric zero-point) and observations of variables near the core of M 15 suffer from photometric contamination. The calibrated LMC data is subsequently added to the Wesenheit template since that galaxy exhibits precise OGLE photometry for innumerable variables of differing classes, that includes recent observations for δ Scuti variables indicating the stars follow a steeper V I Wesenheit function than classical Cepheids pulsating in the fundamental mode. V I photometry for the calibrators is tabulated to facilitate further research, and includes new observations acquired via the AAVSO’s robotic telescope network (e.g., VY Pyx: ⟨V ⟩ = 7.25 and ⟨V ⟩ − ⟨I⟩ = 0.67). The approach outlined here supersedes the lead author’s prior first-order effort to unify variables of the instability strip in order to establish reliable distances.

Accepted for publication in JAAVSO Available from arXiv:1007.2300

Massive runaway stars in the Small Magellanic Cloud

V.V. Gvaramadze^{1,2,3}, J. Pflamm-Altenburg^{1} and P. Kroupa^{1}

1Argelander-Institut für Astronomie, Universität Bonn, Auf dem Hügel 71, 53121 Bonn, Germany 2Sternberg Astronomical Institute, Moscow State University, Universitetskij Pr. 13, Moscow 119992, Russia 3Isaac Newton Institute of Chile, Moscow Branch, Universitetskij Pr. 13, Moscow 119992, Russia

Using archival Spitzer Space Telescope data, we identified for the first time a dozen runaway OB stars in the Small Magellanic Cloud (SMC) through the detection of their bow shocks. The geometry of detected bow shocks allows us to infer the direction of motion of the associated stars and to determine their possible parent clusters and associations. One of the identified runaway stars, AzV 471, was already known as a high-velocity star on the basis of its high peculiar radial velocity, which is offset by ≃ 40 km s^{−1} from the local systemic velocity. We discuss implications of our findings for the problem of the origin of field OB stars. Several of the bow shock-producing stars are found in the confines of associations, suggesting that these may be “alien” stars contributing to the age spread observed for some young stellar systems. We also report the discovery of a kidney-shaped nebula attached to the early WN-type star SMC-WR3 (AzV 60a). We interpreted this nebula as an interstellar structure created owing to the interaction between the stellar wind and the ambient interstellar medium.

Accepted for publication in MNRAS Available from arXiv:1010.2490

The masses, and the mass discrepancy of O-type stars

Carsten Weidner^{1} and Jorick S. Vink^{2}

1Scottish Universities Physics Alliance (SUPA), School of Physics and Astronomy, University of St. Andrews, North Haugh, St. Andrews, Fife KY16 9SS, UK 2Armagh Observatory, College Hill, Armagh BT61 9DG, UK

Context. The ”mass discrepancy” in massive O stars represents a long-standing problem in stellar astrophysics with far-reaching implications for the chemical and dynamical feedback in galaxies. Aims. Our goal is to investigate this mass discrepancy by comparing state-of-the-art model masses with modelindependent masses determined from eclipsing binaries. Methods. Using stellar evolution models and a recent calibration of stellar parameters for O-star spectral sub-classes, we present a convenient way to convert observed solar metallicity O star spectral types into model masses, which we subsequently compare to our dynamical mass compilation. We also derive similar conversions for Large and Small Magellanic Cloud metallicities. Results. We obtain a good agreement between model and dynamical masses, suggesting the long-standing problem of a systematic mass discrepancy problem may have been solved.We also provide error ranges for the model masses, as well as minimal and maximal age estimates for when the model stars are in a given spectral type box.

Accepted for publication in A&A Available from arXiv:1010.2204

Rotation of the Milky Way and the formation of the Magellanic Stream

Adam Růžička^{1,2}, Christian Theis^{1,3} and Jan Palouš^{2}

1Institut für Astronomie der Universität Wien, Austria 2Astronomical Institute of the Academy of Sciences of the Czech Republic, Prague 3Planetarium Mannheim, Germany

We studied the impact of the revisited values for the LSR circular velocity of the Milky Way (Reid et al. 2004) on the formation of the Magellanic Stream. The LSR circular velocity was varied within its observational uncertainties as a free parameter of the interaction between the Large (LMC) and the Small (SMC) Magellanic Clouds and the Galaxy. We have shown that the large-scale morphology and kinematics of the Magellanic Stream may be reproduced as tidal features, assuming the recent values of the proper motions of the Magellanic Clouds (Kallivayalil et al. 2006). Automated exploration of the entire parameter space for the interaction was performed to identify all parameter combinations that allow for modeling the Magellanic Stream. Satisfactory models exist for the dynamical mass of the Milky Way within a wide range of 0.6 × 10^{12} M_{⊙} to 3.0 × 10^{12} M_{⊙} and over the entire 1-σ errors of the proper motions of the Clouds. However, the successful models share a common interaction scenario. The Magellanic Clouds are satellites of the Milky Way, and in all cases two close LMC–SMC encounters occurred within the last 4 Gyr at t < −2.5 Gyr and t ≈ −150 Myr, triggering the formation of the Stream and of the Magellanic Bridge, respectively. The latter encounter is encoded in the observed proper motions and inevitable in any model of the interaction. We conclude that the tidal origin of the Magellanic Stream implies the previously introduced LMC/SMC orbital history, unless the parameters of the interaction are revised substantially.

Accepted for publication in Astrophysical Journal Available from arXiv:1010.0942

Orbital period determinations for four SMC Be/X-ray binaries

M.P.E. Schurch^{1}, M.J. Coe^{2}, V.A. McBride^{2}, L.J. Townsend^{2}, A. Udalski^{3}, F. Haberl^{4} and R.H.D. Corbet^{5}

1Astrophysics, Cosmology and Gravity Centre (ACGC), Astronomy Department, University of Cape Town, Rondebosch, Private Bag X1, 7701, South Africa 2School of Physics and Astronomy, Southampton University, Highfield, Southampton, SO17 1BJ, UK 3Warsaw University Observatory, Aleje Ujazdowskie 4, 00-478 Warsaw,Poland 4Max-Planck-Institut für extraterrestrische Physik, Gießenbachstraße, 85748, Germany 5University of Maryland Baltimore County, X-ray Astrophysics Laboratory, Mail Code 662, NASA Goddard Space Flight Center, Greenbelt, MD 20771, USA

We present an optical and X-ray study of four Be/X-ray binaries located in the Small Magellanic Cloud (SMC). OGLE I-band data of up to 11 years of semi-continuous monitoring has been analysed for SMC X-2, SXP 172 and SXP 202B, providing both a measurement of the orbital period (P_{orb} = 18.62, 68.90, and 229.9 days for the pulsars respectively) and a detailed optical orbital profile for each pulsar. For SXP 172 this has allowed a direct comparison of the optical and X-ray emission seen through regular RXTE monitoring, revealing that the X-ray outbursts precede the optical by around 7 days. Recent X-ray studies by XMM–Newton have identified a new source in the vicinity of SXP 15.3 raising doubt on the identification of the optical counterpart to this X-ray pulsar. Here we present a discussion of the observations that led to the proposal of the original counterpart and a detailed optical analysis of the counterpart to the new X-ray source, identifying a 21.7 d periodicity in the OGLE I-band data. The optical characteristics of this star are consistent with that of a SMC Be/X-ray binary. However, this star was rejected as the counterpart to SXP 15.3 in previous studies due to the lack of Hα emission.

Accepted for publication in Monthly Notices of the Royal Astronomical Society Available from arXiv:1010.5621

Deuterated species in extragalactic star-forming regions

E. Bayet^{1}, Z. Awad^{1} and S. Viti^{1}

1Department of Physics and Astronomy, UCL, UK

We present a theoretical study of the deuterated species detectability in various types of extragalactic star-forming regions based on our predictions of chemical abundances. This work is motivated by the past and current attempts at observing deuterated species in external galaxies such as NGC 253, IC 342 and the LMC. Here, we investigate the influence of the density, the temperature, the FUV radiation field, the cosmic ray ionisation, and the metallicity on the fractional abundances and D/H abundance ratios of about 20 deuterated species. Without modelling any particular source, we determined how the deuterium chemistry behaves in different physical environments such as starburst, cosmic-rays enhanced environments, low metallicity and high redshift galaxies. In general, our predicted column densities seem in good agreement with those derived from the current limited dataset of observations in external galaxies. We provide, for the first time, a list of key deuterated species whose abundances are high enough to be possibly detectable by the Atacama Large Millimeter Array (ALMA) and Herschel, as a function of galactic nuclear activity and redshift.

Published in ApJ Available from arXiv:1010.2868 and from http://www.lra.ens.fr/∼bayet/

Dynamics of the Magellanic Clouds in a ΛCDM universe

Michael Boylan-Kolchin^{1,2}, Gurtina Besla^{3} and Lars Hernquist^{3}

1Max Planck Institute for Astrophysics, Germany 2Center for Galaxy Evolution, UC Irvine, USA 3Harvard-Smithsonian Center for Astrophysics, USA

We examine Milky Way–Magellanic Cloud systems selected from the Millennium-II Simulation in order to place the orbits of the Magellanic Clouds in a cosmological context. Our analysis shows that satellites massive enough to be LMC analogs are typically accreted at late times. Moreover, those that are accreted at early times and survive to the present have orbital properties that are discrepant with those observed for the LMC. The high velocity of the LMC, coupled with the dearth of unbound orbits seen in the simulation, argues that the mass of the MW’s halo is unlikely to be less than 2 × 10^{12} M_{⊙}. This conclusion is further supported by statistics of halos hosting satellites with masses, velocities, and separations comparable to those of the LMC. We further show that: (1) LMC and SMC-mass objects are not particularly uncommon in MW-mass halos; (2) the apparently high angular momentum of the LMC is not cosmologically unusual; and (3) it is rare for a MW halo to host a LMC–SMC binary system at z = 0, but high speed binary pairs accreted at late times are possible. Based on these results, we conclude that the LMC was accreted within the past four Gyr and is currently making its first pericentric passage about the MW.

Submitted to MNRAS Available from arXiv:1010.4797

Could the Magellanic Clouds be tidal dwarves expelled from a past-merger event occurring in Andromeda?

Yanbin Yang^{1,2} and François Hammer^{2}

1National Astronomical Observatories, Chinese Academy of Sciences, 20A Datun Road, Chaoyang District, Beijing 100012, China 2GEPI, Observatoire de Paris, CNRS, Université Paris Diderot; 5 Place Jules Janssen, Meudon, France

The Magellanic Clouds are often considered as outliers in the satellite system of the Milky Way because they are irregular and gas-rich galaxies. From their large relative motion, they are likely from their first pass near the Milky Way, possibly originating from another region of the Local Group or its outskirts. M 31 could have been in a merger stage in its past and we investigate whether or not the Large Magellanic Cloud could have been a tidal dwarf expelled during this event. Such an hypothesis is tested in the frame of present-day measurements and uncertainties of the relative motions of LMC and M 31. Our method is to trace back the LMC trajectory using several thousands of different configurations that sample the corresponding parameter space. We find several configurations that let LMC at 50 kpc from M 31, 4.3 to 8 Gyrs ago , depending on the adopted shape of the Milky Way halo. For all configurations, the LMC velocity at such a location is invariably slightly larger than the escape velocity at such a radius. The preferred solutions correspond to a spherical to prolate Milky Way halo, predicting a transversal motion of M 31 of less than 107 km s^{−1} and down to values that are close to zero. We conclude that from present-day measurements, Magellanic Clouds could well be tidal dwarves expelled from a former merger events occurring in M 31.

Accepted for publication in The Astrophysical Journal Letters Available from arXiv:1010.2748

Searching for Galactic hidden gas through interstellar scintillation: Results from a test with the NTT–SOFI detector

Farhang Habibi^{1,2}, Marc Moniez^{1}, Reza Ansari^{1} and Sohrab Rahvar^{1,2}

1Laboratoire de l’Accélérateur Linéaire, IN2P3-CNRS, Université de Paris-Sud, B.P. 34, 91898 Orsay Cedex, France 2Department of Physics, Sharif University of Technology P.O. Box 11365-9161, Tehran, Iran

Aims: Stars twinkle because their light propagates through the atmosphere. The same phenomenon is expected at a longer time scale when the light of remote stars crosses an interstellar molecular cloud, but it has never been observed at optical wavelength. In a favorable case, the light of a background star can be subject to stochastic fluctuations on the order of a few percent at a characteristic time scale of a few minutes. Our ultimate aim is to discover or exclude these scintillation effects to estimate the contribution of molecular hydrogen to the Galactic baryonic hidden mass. This feasibility study is a pathfinder toward an observational strategy to search for scintillation, probing the sensitivity of future surveys and estimating the background level. Methods: We searched for scintillation induced by molecular gas in visible dark nebulae as well as by hypothetical halo clumpuscules of cool molecular hydrogen (H_{2}–He) during two nights. We took long series of 10s infrared exposures with the ESO–NTT telescope toward stellar populations located behind visible nebulae and toward the Small Magellanic Cloud (SMC). We therefore searched for stars exhibiting stochastic flux variations similar to what is expected from the scintillation effect. According to our simulations of the scintillation process, this search should allow one to detect (stochastic) transverse gradients of column density in cool Galactic molecular clouds of order of ∼ 3 × 10^{−5} g cm^{−2}/10 000 km. Results: We found one light-curve that is compatible with a strong scintillation effect through a turbulent structure characterized by a diffusion radius R_{diff} < 100 km in the B 68 nebula. Complementary observations are needed to clarify the status of this candidate, and no firm conclusion can be established from this single observation. We can also infer limits on the existence of turbulent dense cores (of number density n > 10^{9} cm^{−3}) within the dark nebulae. Because no candidate is found toward the SMC, we are also able to establish upper limits on the contribution of gas clumpuscules to the Galactic halo mass. Conclusions: The limits set by this test do not seriously constrain the known models, but we show that the short time-scale monitoring for a few 10^{6} star × hour in the visible band with a > 4 meter telescope and a fast readout camera should allow one to quantify the contribution of turbulent molecular gas to the Galactic halo. The LSST (Large Synoptic Survey Telescope) is perfectly suited for this search.

Accepted for publication in Astronomy and Astrophysics Available from arXiv:1010.4616

The XMM–Newton survey of the Small Magellanic Cloud: Discovery of

the 11.866 s Be/X-ray binary pulsar XMMU J004814.0−732204 (SXP 11.87)

R. Sturm^{1}, F. Haberl^{1}, M.J. Coe^{2}, E.S. Bartlett^{2}, D.A.H. Buckley^{3}, R.H.D. Corbet^{4}, M. Ehle^{5}, M.D. Filipović^{6}, D. Hatzidimitriou^{7,8}, S. Mereghetti^{9}, N. La Palombara^{9}, W. Pietsch^{1}, A. Tiengo^{9}, L.J. Townsend^{2} and A. Udalski^{10}

1Max-Planck-Institut für extraterrestrische Physik, Gießenbachstraße, 85748 Garching, Germany 2School of Physics and Astronomy, University of Southampton, Highfield, Southampton SO17 1BJ, United Kingdom 3South African Astronomical Observatory, P.O. Box 9, Observatory 7935, Cape Town, South Africa 4University of Maryland, Baltimore County, Mail Code 662, NASA Goddard Space Flight Center, Greenbelt, MD 20771, USA 5XMM–Newton Science Operations Centre, ESAC, ESA, PO Box 78, 28691 Villanueva de la Cañada, Madrid, Spain 6University of Western Sydney, Locked Bag 1797, Penrith South DC, NSW1797, Australia 7Department of Astrophysics, Astronomy and Mechanics, Faculty of Physics, University of Athens, Panepistimiopolis, GR15784 Zografos, Athens, Greece 8Foundation for Research and Technology Hellas, IESL, Greece 9INAF, Istituto di Astrofisica Spaziale e Fisica Cosmica Milano, via E. Bassini 15, 20133 Milano, Italy 10Warsaw University Observatory, Aleje Ujazdowskie 4, 00-478 Warsaw, Poland

One of the goals of the XMM–Newton survey of the Small Magellanic Cloud is the study of the Be/X-ray binary population. During one of our first survey observations a bright new transient – XMMU J004814.0−732204 – was discovered. We present the analysis of the EPIC X-ray data together with optical observations, to investigate the spectral and temporal characteristics of XMMU J004814.0−732204. We found coherent X-ray pulsations in the EPIC data with a period of (11.86642 ± 0.00017) s. The X-ray spectrum can be modelled by an absorbed power-law with indication for a soft excess. Depending on the modelling of the soft X-ray spectrum, the photon index ranges between 0.53 and 0.66. We identify the optical counterpart as a B = 14.9 mag star which was monitored during the MACHO and OGLE-III projects. The optical light curves show regular outbursts by ∼0.5 mag in B and R and up to 0.9 mag in I which repeat with a time scale of about 1000 days. The OGLE-III optical colours of the star are consistent with an early B spectral type. An optical spectrum obtained at the 1.9-m telescope of the South African Astronomical Observatory in December 2009 shows Hα emission with an equivalent width of 3.5 ± 0.6 Å. The X-ray spectrum and the detection of pulsations suggest that XMMU J004814.0−732204 is a new high mass X-ray binary pulsar in the SMC. The long term variability and the Hα emission line in the spectrum of the optical counterpart identify it as a Be/X-ray binary system.

Accepted for publication in Astronomy and Astrophysics Available from arXiv:1010.5695

How common are the Magellanic Clouds

Lulu Liu^{1}, Brian F. Gerke^{1}, Risa H. Wechsler^{1}, Peter S. Behroozi^{1} and Michael T. Busha^{1}

1KIPAC, Stanford University, SLAC National Accelerator Laboratory, USA

We introduce a probabilistic approach to the problem of counting dwarf satellites around host galaxies in databases with limited redshift information. This technique is used to investigate the occurrence of satellites with luminosities similar to the Magellanic Clouds around hosts with properties similar to the Milky Way in the object catalog of the Sloan Digital Sky Survey. Our analysis uses data from SDSS Data Release 7, selecting candidate Milky-Way-like hosts from the spectroscopic catalog and candidate analogs of the Magellanic Clouds from the photometric catalog. Our principal result is the probability for a Milky-Way-like galaxy to host N_{sat} close satellites with luminosities similar to the Magellanic Clouds. We find that 81 percent of galaxies like the Milky Way are have no such satellites within a radius of 150 kpc, 11 percent have one, and only 3.5 percent of hosts have two. The probabilities are robust to changes in host and satellite selection criteria, background-estimation technique, and survey depth. These results demonstrate that the Milky Way has significantly more satellites than a typical galaxy of its luminosity; this fact is useful for understanding the larger cosmological context of our home galaxy.

Submitted to ApJ Available from arXiv:1011.2255

Ice chemistry in massive Young Stellar Objects: the role of metallicity

J.M. Oliveira^{1}, J.Th. van Loon^{1}, G.C. Sloan^{2}, R. Indebetouw^{3,4}, F. Kemper^{5}, A.G.G.M. Tielens^{6}, J.D. Simon^{7}, Paul M. Woods^{5} and M. Meixner^{8}

1Lennard-Jones Laboratories, School of Physical and Geographical Sciences, Keele University, Staffordshire ST5 5BG, UK 2Department of Astronomy, Cornell University, Ithaca, NY 14853, USA 3Department of Astronomy, University of Virginia, P.O. Box 400325, Charlottesville, VA 22904, USA 4National Radio Astronomy Observatory, 520 Edgemont Road, Charlottesville, VA 22903, USA 5Jodrell Bank Centre for Astrophysics, School of Physics and Astronomy, The University of Manchester, Manchester M13 9PL, UK 6Leiden Observatory, P.O. Box 9513, NL-2300 RA Leiden, The Netherlands 7Observatories of the Carnegie Institution of Washington, 813 Santa Barbara St., Pasadena, CA 91101, USA 8Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA

We present the comparison of the three most important ice constituents (water, CO and CO_{2}) in the envelopes of massive Young Stellar Objects (YSOs), in environments of different metallicities: the Galaxy, the Large Magellanic Cloud (LMC) and, for the first time, the Small Magellanic Cloud (SMC). We present observations of water, CO and CO_{2} ice in 4 SMC and 3 LMC YSOs (obtained with Spitzer-IRS and VLT/ISAAC). While water and CO_{2} ice are detected in all Magellanic YSOs, CO ice is not detected in the SMC objects. Both CO and CO_{2} ice abundances are enhanced in the LMC when compared to high-luminosity Galactic YSOs. Based on the fact that both species appear to be enhanced in a consistent way, this effect is unlikely to be the result of enhanced CO_{2} production in hotter YSO envelopes as previously thought. Instead we propose that this results from a reduced water column density in the envelopes of LMC YSOs, a direct consequence of both the stronger UV radiation field and the reduced dust-to-gas ratio at lower metallicity. In the SMC the environmental conditions are harsher, and we observe a reduction in CO_{2} column density. Furthermore, the low gas-phase CO density and higher dust temperature in YSO envelopes in the SMC seem to inhibit CO freeze-out. The scenario we propose can be tested with further observations.

Accepted for publication in MNRAS Letters Available from arXiv:1011.2786

Survey of O vi in the LMC

Amit Pathak^{1}, A.C. Pradhan^{2}, N.V. Sujatha^{3} and Jayant Murthy^{4}

1Indian Institute of Astrophysics, India

We present a survey of interstellar O vi absorption in the Large Magellanic Cloud (LMC) towards 70 lines of sight based on Far Ultraviolet Spectroscopic Explorer (FUSE) observations. The survey covers O vi absorption in a large number of objects in different environmental conditions of the LMC. Overall, a high abundance of O vi is present in active and inactive regions of the LMC with mean log N(OVI) = 14.23 atoms cm^{−2}. There is no correlation observed between O vi absorption and emissions from the hot gas (X-ray surface brightness) or the warm gas (Hα surface brightness). O vi absorption in the LMC is patchy and the properties are similar to that of the Milky Way (MW). In comparison to the Small Magellanic Cloud (SMC), O vi is lower in abundance even though SMC has a lower metallicity compared to the LMC and the MW. We present observations in 10 superbubbles of the LMC of which we detect O vi absorption in 5 superbubbles for the first time and the superbubbles show an excess O vi absorption of about 40% compared to non-superbubble lines of sight. We have also studied the properties of O vi absorption in the 30 Doradus region. Even though O vi does not show any correlation with X-ray emission for the LMC, a good correlation between log N(OVI) and X-ray surface brightness for 30 Doradus region is present. We also find that O vi abundance decreases with increasing distance from the star cluster R 136.

Accepted for publication in MNRAS Available from arXiv:1011.0690

Lightcurve solutions of eclipsing binaries in the SMC

Valentin P. Ivanov^{1}, Diana P. Kjurkchieva^{1} and M. Srinivasa Rao^{2}

1Department of Physics, Shumen University, 9700 Shumen, Bulgaria 2Indian Institute of Astrophysics, Bangalore 560034, India

We propose a procedure for light-curve solution of eclipsing binary stars in the Small Magellanic Cloud for which photometric data have been obtained in the framework of the OGLE project as well as way of determination of the global stellar parameters on the basis of the obtained solutions, some empirical relations as well as the distance to the SMC. Several examples illustrate this procedure.

Accepted for publication in Bulletin of the Astronomical Society of India (BASI) Available from arXiv:1011.1986 and from http://www.ncra.tifr.res.in/∼basi/accepted.htm

New limb-darkening coefficients and synthetic photometry for model-atmosphere grids at Galactic, LMC, and SMC abundances.

Ian D. Howarth^{1}

1Dept. Physics & Astronomy, UCL, Gower St., London WC1E 6BT, UK

New grids of Atlas9 models have been calculated using revised convection parameters and updated opacity-distribution functions,for chemical compositions intended to be representative of solar, [M/H] = +0.3, +0.5, Large Magellanic Cloud (LMC), and Small Magellanic Cloud (SMC) abundances. The grids cover T_{eff} = 3.5–50 kK, from log g = 5.0 to the effective Eddington limit. Limb-darkening coefficients and synthetic photometry are presented in the UBV RIJHKLM, uvby, ugriz, WFCAM, Hipparcos/Tycho, and Kepler passbands for these models, and for Castelli’s comparable ‘new-ODF’ grids. Flux distributions are given for the new models. The sensitivity of limb-darkening coefficients to the adopted physics is illustrated.

Submitted to MNRAS Available from arXiv:1011.2631 and from http://zuserver2.star.ucl.ac.uk/∼idh/NewGrids/

Data characterization using artificial-star tests: performance evaluation

Yi Hu^{1}, Licai Deng^{1}, Richard de Grijs^{2} and Qiang Liu^{1}

1National Astronomical Observatories, Chinese Academy of Sciences, Chine 2Kavli Institute for Astronomy and Astrophysics, Peking University, China

Traditional artificial-star tests are widely applied to photometry in crowded stellar fields. However, to obtain reliable binary fractions (and their uncertainties) of remote, dense, and rich star clusters, one needs to recover huge numbers of artificial stars. Hence, this will consume much computation time for data reduction of the images to which the artificial stars must be added. In this paper, we present a new method applicable to data sets characterized by stable, well-defined point-spread functions, in which we add artificial stars to the retrieved-data catalog instead of the raw images. Taking the young Large Magellanic Cloud cluster NGC 1818 as an example, we compare results from both methods and show that they are equivalent, while our new method saves significant computational time.

Accepted for publication in PASP Available from arXiv:1011.3892

The mass distribution and assembly history of the Milky Way from the properties of the Magellanic Clouds

Michael T. Busha^{1,2}, Philip J. Marshall^{1,3}, Risa H. Wechsler^{1,4}, Anatoly Klypin^{5} and Joel Primack^{6}

1Stanford University, USA 2University of Zurich, Switzerland 3Oxford University, UK 4SLAC National Accelerator Laboratory, USA 5New Mexico State University, USA 6University of California Santa Cruz, USA

We present a new measurement of the mass of the Milky Way (MW) based on observed properties of its largest satellite galaxies, the Magellanic Clouds (MCs), and an assumed prior of a ΛCDM universe. A large, high resolution cosmological simulation of this universe provides a means to statistically sample the dynamical properties of bright satellite galaxies in a large population of dark matter halos. The observed properties of the MCs, including their circular velocity, distance from the center of the MW, and velocity within the MW halo, are used to evaluate the likelihood that a given halo would have each or all of these properties; the posterior PDF for any property of the MW system can thus be constructed. This method provides a constraint on the MW virial mass, 1.2_{−0.4}^{+0.7} × 10^{12} M_{⊙} (68% confidence), which is consistent with recent determinations that involve very different assumptions. In addition, we calculate the posterior PDF for the density profile of the MW and its satellite accretion history. Although typical satellites of 10^{12} M_{⊙} halos are accreted over a wide range of epochs over the last 10 Gyr, we find a 72% probability that the Magellanic Clouds were accreted within the last Gyr, and a 50% probability that they were accreted together.

Submitted to Astrophysical Journal Available from arXiv:1011.2203

Period–luminosity relations of type II Cepheids in the Magellanic Clouds

Noriyuki Matsunaga^{1}, Michael Feast^{2,3} and Igor Soszyński^{4}

1Kiso Observatory, University of Tokyo, 10762-30, Mitake, Kiso, Nagano 397-0101, Japan 2University of Cape Town, Rondebosch, 7701, South Africa 3South African Astronomical Observatory, P.O. Box 9, Observatory, 7935, South Africa 4Warsaw University, Al. Ujazdowskie 4, 00-478 Warszawa, Poland

Period–luminosity relations (PLRs) of type II Cepheids (T2Cs) in the Small Magellanic Cloud are derived based on OGLE-III, IRSF/SIRIUS and other data, and these are compared with results for the Large Magellanic Cloud and Galactic globular clusters. Evidence is found for a change of the PLR slopes from system to system. Treating the longer period T2Cs (W Vir stars) eparately gives an SMC–LMC modulus difference of 0.39 ± 0.05 mag without any metallicity corrections being applied. This agrees well with the difference in moduli based on different distance indicators, in particular the PLRs of classical Cepheids. The shorter period T2Cs (BL Her stars) give a smaller SMC– LMC difference suggesting that their absolute magnitudes might be affected either by metallicity or by age effects. It is shown that the frequency distribution of T2C periods also changes from system to system.

Accepted for publication in MNRAS Available from arXiv:1012.0098

Period–magnitude relation of Mira-like variables in the Large Magellanic Cloud as a tool to understand circumstellar extinction

Yoshifusa Ita^{1} and Noriyuki Matsunaga^{2}

1Astronomical Institute, Graduate School of Science, Tohoku University, 6-3 Aramaki Aoba, Aoba-ku, Sendai, Miyagi 980-8578, Japan 2Institute of Astronomy, School of Science, The University of Tokyo, Mitaka, Tokyo 181-0015, Japan

Near- to mid-infrared period–magnitude relations and also the period–bolometric luminosity relation of OGLE-III Mira-like variables in the LMC are derived. The relations have a kink, and the period at which the break occurs is quantitatively obtained. There are many Mira-like variables whose fluxes at the optical and the near-infrared wavebands are fainter than the ones predicted by the period–magnitude relations. The deviation is due to the circumstellar extinction, and the amount of the deviation is found to be strongly correlated with near-infrared colors. The empirical formulae relating the amount of the deviation and the near-infrared colors are derived. These relations are useful to accurately calculate the distances to the dusty Mira-like variables, because the dimmed fluxes due to the circumstellar extinction can be estimated. In a manner analogous to the interstellar extinction law, the ratios of deviations at any two different wavebands are calculated. The ratios are found to change with the pulsation period, indicating that the dust properties are subject to change as Mira-like variables evolve.

Accepted for publication in MNRAS Available from arXiv:1011.4984

Conference Papers

Infrared properties of Active OB stars in the Magellanic Clouds from the Spitzer SAGE Survey

A.Z. Bonanos^{1}, D.J. Lennon^{2}, D.L. Massa^{2}, M. Sewilo^{2}, F. Köhlinger^{3}, N. Panagia^{2}, J.Th. van Loon^{4}, C.J. Evans^{5}, L.J. Smith^{2}, M. Meixner^{2} and K. Gordon^{2}

1Institute of Astronomy & Astrophysics, National Observatory of Athens, I. Metaxa & Vas. Pavlou St., P. Penteli, 15236 Athens, Greece 2Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 3Department of Physics and Astronomy, Heidelberg University, Albert-Überle-Str. 3-5, D-69120 Heidelberg, Germany 4Astrophysics Group, Lennard-Jones Laboratories, Keele University, Staffordshire ST5 5BG, UK 5UK Astronomy Technology Centre, Royal Observatory Edinburgh, Blackford Hill, Edinburgh, EH9 3HJ, UK

We present a study of the infrared properties of 4922 spectroscopically confirmed massive stars in the Large and Small Magellanic Clouds, focusing on the active OB star population. Besides OB stars, our sample includes yellow and red supergiants, Wolf–Rayet stars, Luminous Blue Variables (LBVs) and supergiant B[e] stars. We detect a distinct Be star sequence, displaced to the red, and find a higher fraction of Oe and Be stars among O and early-B stars in the SMC, respectively, when compared to the LMC, and that the SMC Be stars occur at higher luminosities. We also find photometric variability among the active OB population and evidence for transitions of Be stars to B stars and vice versa. We furthermore confirm the presence of dust around all the supergiant B[e] stars in our sample, finding the shape of their spectral energy distributions (SEDs) to be very similar, in contrast to the variety of SED shapes among the spectrally variable LBVs.

Oral contribution, published in IAUS 272 on ”Active OB stars: structure, evolution, mass loss and critical limits”, Paris, July 19–23, 2010 Available from arXiv:1010.0139

The infrared properties of massive stars in the Magellanic Clouds

A.Z. Bonanos^{1}, D.J. Lennon^{2}, D.L. Massa^{2}, M. Sewilo^{2}, F. Köhlinger^{3}, N. Panagia^{2}, J.Th. van Loon^{4}, C.J. Evans^{5}, L.J. Smith^{2}, M. Meixner^{2}, K. Gordon^{2} and SAGE teams^{2}

1Institute of Astronomy & Astrophysics, National Observatory of Athens, I. Metaxa & Vas. Pavlou St., P. Penteli, 15236 Athens, Greece 2Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 3Department of Physics and Astronomy, Heidelberg University, Albert-Überle-Str. 3-5, D-69120 Heidelberg, Germany 4Astrophysics Group, Lennard-Jones Laboratories, Keele University, Staffordshire ST5 5BG, UK 5UK Astronomy Technology Centre, Royal Observatory Edinburgh, Blackford Hill, Edinburgh, EH9 3HJ, UK

We present results of our study of the infrared properties of massive stars in the Large and Small Magellanic Clouds, which are based on the Spitzer SAGE surveys of these galaxies. We have compiled catalogs of spectroscopically confirmed massive stars in each galaxy, as well as photometric catalogs for a subset of these stars that have infrared counterparts in the SAGE database, with uniform photometry from 0.3 to 24 μm in the UBVIJHK_{s}+IRAC+MIPS24 bands. These catalogs enable a comparative study of infrared excesses of OB stars, classical Be stars, yellow and red supergiants, Wolf–Rayet stars, Luminous Blue Variables and supergiant B[e] stars, as a function of metallicity, and provide the first roadmaps for interpreting luminous, massive, resolved stellar populations in nearby galaxies at infrared wavelengths.

Poster contribution, published in 39^{th} Liège Astrophysical Colloquium, ”The multi-wavelength view of hot, massive stars” Available from arXiv:1010.0121

The B[e] phenomenon in the Milky Way and Magellanic Clouds

Anatoly Miroshnichenko^{1}, Nadine Manset^{2}, Francesco Polcaro^{3}, Corinne Rossi^{4} and Sergey Zharikov^{5}

1University of North Carolina at Greensboro, Greensboro, NC, USA 2Canada–France–Hawai’i Telescope Corporation, Kamuela, HI, USA 3Instituto di Astrofisica Spaziale e Fisica Cosmica, INAF, Roma, Italy 4Università La Sapienza Roma, Italy 5Instituto de Astronomía, Universidad Nacional Autónoma de México, Ensenada, México

Discovered over 30 years ago, the B[e] phenomenon has not yet revealed all its puzzles. New objects that exhibit it are being discovered in the Milky Way, and properties of known objects are being constrained. We review recent findings about objects of this class and their subgroups as well as discuss new results from studies of the objects with yet unknown nature. In the Magellanic Clouds, the population of such objects has been restricted to supergiants. We present new candidates with apparently lower luminosities found in the LMC.

Oral contribution, published in IAU Symposium 272, ”Active OB stars: structure, evolution, mass loss and critical limits” Available from arXiv:1010.0456

Review Papers

Populations of OB-type stars in galaxies

C.J. Evans^{1}

1UK Astronomy Technology Centre, Edinburgh, UK

One of the challenges for stellar astrophysics is to reach the point at which we can undertake reliable spectral synthesis of unresolved populations in young, star-forming galaxies at high redshift. Here I summarise recent studies of massive stars in the Galaxy and Magellanic Clouds, which span a range of metallicities commensurate with those in highredshift systems, thus providing an excellent laboratory in which to study the role of environment on stellar evolution. I also give an overview of observations of luminous supergiants in external galaxies out to a remarkable 6.7 Mpc, in which we can exploit our understanding of stellar evolution to study the chemistry and dynamics of the host systems.

Published in IAU Symposium 272, eds. Neiner, Wade, Meynet & Peters (invited review) Available from arXiv:1009.4484

Massive stellar models: rotational evolution, metallicity effects

Sylvia Ekström^{1}, Cyril Georgy^{1}, Georges Meynet^{1}, André Maeder^{1} and Anahí Granada^{1,2}

1Geneva Observatory, University of Geneva, Switzerland 2Instituto de Astrofísica de La Plata, Universidad Nacional de La Plata, Argentina

The Be star phenomenon is related to fast rotation, although the cause of this fast rotation is not yet clearly established. The basic effects of fast rotation on the stellar structure are reviewed: oblateness, mixing, anisotropic winds. The processes governing the evolution of the equatorial velocity of a single star (transport mechanisms and mass loss) are presented, as well as their metallicity dependence. The theoretical results are compared to observations of B and Be stars in the Galaxy and the Magellanic Clouds.

Published in IAU Symposium 272, ”Active OB stars – structure, evolution, mass loss, and critical limits”, Paris, France, July 19–23, 2010 Available from arXiv:1010.3838

Job Adverts

Postdoctoral position in star cluster evolution and resolved stellar populations

Applications are invited for a postdoctoral position based at the Kavli Institute for Astronomy and Astrophysics (KIAA) at Peking University. The successful applicant will work with Prof. Richard de Grijs on a programme of forefront research into the formation and evolution of star clusters and their galactic-field environments. Depending on the successful applicant’s background and expertise, the project will involve either (i) analysis of high-resolution spectroscopy combined with aspects of numerical modelling or (ii) resolved stellar population studies of the Magellanic Clouds or the Galactic Centre region based on new near-infrared VISTA observations. Applicants are encouraged to propose programmes of their own interest within these constraints.

The Institute is designed to be a forum for global scientific exchange and a training centre for international postdocs and students. KIAA organises a variety of academic activities and programmes to stimulate research and promote interdisciplinary interactions. The working language is English and there is a broad mix of international faculty and postdocs. The successful applicant will be given the opportunity to supervise some of the highly talented Peking University graduate students.

The appointment is grant supported and will be made for two years from 1 March 2011 or as soon as possible thereafter. Informal enquiries are welcomed by Prof. Richard de Grijs (grijs@pku.edu.cn). To apply, send a cover letter, CV, and research plan (maximum 3 pages in 12pt font) to grijs@pku.edu.cn. Assessment of complete applications will begin on 1 January 2011 until a suitable candidate has been identified. Recommendation letters will be taken up for short-listed applicants only.

See also http://kiaa.pku.edu.cn

HERMES Super Science Fellowships

The Department of Physics and Astronomy at Macquarie University seeks to appoint THREE (3) highly-motivated Super Science Fellows to carry out research related to the HERMES project. HERMES is an innovative new multi-object high-resolution spectrograph being built for the Anglo–Australian Telescope, specifically designed for conducting large stellar surveys, which will observe more than 1 million stars in the Galactic disk and Magellanic Clouds in order to test current theories of galaxy formation and stellar astrophysics (see http://www.aao.gov.au/HERMES).

The Super Science Fellows will work with Dr. Daniel Zucker and the other members of the HERMES Super Science team (Freeman, Bland-Hawthorn, Lattanzio, De Silva), as well as with collaborators at the Australian Astronomical Observatory, Australian National University, the University of Sydney, Monash University and other universities and institutes around the world. The Fellows will lay the groundwork for, and exploit the data flowing from, the massive Galactic Archaeology and Magellanic Cloud surveys planned for HERMES. These surveys will obtain detailed elemental abundances and precision radial velocities for over a million stars, yielding a dataset uniquely suited to a wide range of astrophysical research, and directly complementary to major projects such as ESA’s Gaia mission.

Macquarie University, a dynamic research and teaching institution, is located on a park-like campus in suburban Sydney, Australia, in close proximity to the headquarters of the Australian Astronomical Observatory (AAO) and the Australia Telescope National Facility (ATNF). The Astronomy and Astrophysics group at Macquarie has undergone a rapid expansion in the past 5 years, now comprising over 35 faculty, postdoctoral researchers and students, making it one of the largest and most active astronomy groups within Australia. In order to enhance collaboration between the participating institutions, the Fellows will have the opportunity to visit the AAO, ANU, Monash, and the University of Sydney for up to a year.

The Fellowships are full-time, fixed-term 3 year positions, available from 1 July 2011, with research funds of AUD$25,000 per year per fellow attached to the project. There are also two PhD scholarships associated with the project, and Fellows will have the opportunity to co-supervise these students. Applicants should have a PhD in a relevant field (e.g., astronomy or physics), usually awarded not more than 3 years before the nominal start date for the Fellowships (1 July 2011); a demonstrated track record of ongoing research and publication, with an emphasis on observational or theoretical astronomy; the ability to carry out independent research and the ability or capacity to develop research collaborations; the ability or capacity to attract and supervise research students, and to obtain competitive funding; the ability to work independently and as a member of a team; and excellent written, verbal and interpersonal communication skills.

Applications must be submitted through the Macquarie University jobs website: http://www.mq.edu.au/jobs , Reference No. 00V8B.

Candidates should include a CV, list of publications, contact details for three referees, and a one to two page statement about their research interests and how these would contribute to the HERMES project. Applications must be received by 15 December 2010.

The positions have an annual salary package from AUD$91,916 per annum, including (Level B) base salary from AUD$77,670 to AUD$92,103, plus 17% employer’s superannuation contribution and annual leave loading. Relocation assistance will be provided in accordance with the Macquarie University Relocation Guidelines.

See also http://www.aao.gov.au/HERMES (information about HERMES) http://www.mq.edu.au/jobs (to apply for the position)